OpenStudy (anonymous):

Write a polynomial function in standard form with the given zeros x=-1,3,4

OpenStudy (anonymous):

(x+1)(x-3)(x-4)

OpenStudy (anonymous):

given zeros are the solutions so you put in the opposite added to x! then foil it out!

OpenStudy (anonymous):

(x+1)(x-3)= x^2 -2x -3

OpenStudy (anonymous):

x^2 -2x -3 (x-4) = x^3 -2x^2 -3x -4x^2 +8x +12

OpenStudy (anonymous):

simplify = x^3 -6x^2 +5x +12 = y